Tan( pi/3 + x)* tan( pi /3-x) = 2 cos2x +1/ 2 cos2x -1

Grade:11

1 Answers

na batau
5 years ago

tan(A+B).tan(A-B)={sin(A+B).sin(A-B)}/{cos(A+B).cos(A-B)}

use formlas to get

tan(A+B).tan(A-B)={(cos(B))^2-(cos(A))^2}/{(cos(B))^2-(sin(A))^2}

on RHS

use cos2x=2(cos(x))^2-1 to get the LHS